# Inertial Torsion Noise in Matter-Wave Interferometers for Gravity Experiments

### v1

@@ Abstract
Matter-wave interferometry is susceptible to non-inertial noise sources, which can induce dephasing and a resulting loss of interferometric visibility. Here, we focus on inertial torsion noise (ITN), which arises from the rotational motion of the experimental apparatus suspended by a thin wire and subject to random external torques. We provide analytical expressions for the ITN noise starting from Langevin equations describing the experimental box in a thermal environment which can then be used together with the transfer function to obtain the dephasing factor. We verify the theoretical modelling and the validity of the approximations using Monte Carlo simulations obtaining good agreement between theory and numerics. As an application we estimate the size of the effects for the next-generation of interferometery experiments with femtogram particles, which could be used as the building block for entanglement-based tests of the quantum nature of gravity. We find that the ambient gas is a weak source of ITN, posing mild restrictions on the ambient pressure and temperature, and conclude with a discussion about the general ITN constrains by assuming a Langevin equation parameterized by three phenomenological parameters.
@@ I Introduction
Matter-wave interferometry has many salient applications for gravitational physics with devices spanning gravimeters [1], gradiometers [2, 3], accelerometers [4] and gyroscopes [5]. They can also be used for fundamental physics, such as testing the equivalence principle [6, 7, 8] and the quantum gravity induced entangled of masses (QGEM) to test the quantum nature of gravity in a lab [9, 10, 11, 12, 13]  [^1] The first reporting of the testing the quantum nature of gravity via entanglement was done in a symposium [14]. The paper [13] provides the protocol to test the spin-2 nature of the graviton in an analogue of the light-bending experiment, see also [12]. One can also probe the nature of massive graviton [15] and non-local gravitational interaction [16] motivated by string theory.. Furthermore, some groups even consider building gravitational wave observatories based on matter-wave interferometry such as the matter-wave laser interferometer gravitation antenna (MIGA) [17, 18], the matter-wave atomic gradiometer interferometric sensor (MAGIS-100) [19, 20], as well as the mesoscopic interference for metric and curvature (MIMAC) scheme [21].
Future matter-wave interferometry aims to exploit the regime of large masses, large superposition sizes, and long coherence times, allowing the probe of exquisitely small experimental signals. For example, the (QGEM) experiment would ideally require a test mass of $\sim 10^{-15}$ kg, a superposition size of $\sim 100\,\rm\mu m$, and a coherence time of $\sim$1 s [9, 8, 22, 23, 24]. One of the most promising setups towards this kind of experiments is the adaptation of the Stern-Gerlach interferometer (SGI) to nanoparticles [25]. SGIs based on atom chips [26] have already achieved the superposition size and coherence time of 3.93 $\rm\mu m$ and 21.45 ms for the half-loop configuration [27], respectively, and $0.38\,\rm\mu m$ and 7 ms for the full-loop configuration [28], respectively. The next generation of SGIs is currently under theoretical and numerical investigation [29, 30, 31, 32, 33].
@@
This paper will study the dephasing caused by the residual rotational or inertial torsion noise (ITN) for an asymmetric nanoparticle matter-wave interferometer which is sensitive to gravity-gradients [21, 35, 36]. ITN arises naturally in any setup whenever the experimental apparatus is subject to random torques placing it in non-inertial rotational motion. As we will see, ITN can induce relative random phases in an interferometric experiment, resulting in a loss of interferometric contrast. Here, we will be primarily interested in understanding ITN, and we will focus on a simple single-stage suspension forming a torsion pendulum, i.e., matter-wave interferometry performed inside a hanging box. Analogous configurations of the experimental apparatus have been investigated previously in gravitational wave observatories like LIGO [40] and VIRGO [41]. More advanced setups could employ additional structures like the inverted pendulum [42] and the Roberts linkage [43].
In Sec. II, we first introduce the concept of ITN, illustrating it for matter-wave interferometry that can be modelled using qubits. In Sec. III, we generalize the analysis using linear response theory, providing the transfer function and its relation to the dephasing factor. In Sec. IV, we investigate ITN caused by ambient gas collisions on the experimental box starting from the classical Langevin equation. We will compute the power spectrum density (PSD) of the ITN using the convolution theorem and verify the validity of the approximations using Monte Carlo simulations. In Sec. V, we obtain the resulting constraints on the ambient pressure and temperature and general constraints on ITN by parametrizing a generic Langevin equation modelling the experimental box with three phenomenological parameters. In Sec. VII, we conclude with a summary of the obtained results. In Appendix A, we construct the ITN Lagrangian starting from Fermi normal coordinates and transforming to a rotating reference frame. In Appendix B, we provide for completeness also the complete derivation of the ITN PSD using complex analysis.
II Concept of Inertial Torsion Noise
@@
Figure 1: (a) Experimental scheme and illustration of inertial torsion noise (ITN). The matter-wave particle (light and dark blue circles) is placed inside the experimental box (blue box), which is suspended by a thin wire (black vertical line). The box can rotate around the z-axis, while the interferometric protocol is performed along the horizontal x-axis. ITN generates random torques on the experimental box (blue rectangle), placing it in non-inertial rotational motion. The random non-inertial rotational motion induces random relative phases, which can lead to dephasing and the loss of interferometric contrast. (b) The structure of a Stern-Gerlach interferometer with spin-1 [25]. The left path keeps the same position, while the right path accelerates and decelerates at a constant rate resulting in a superposition size $\Delta x$. Such motion is achieved using a constant magnetic field gradient, i.e. $B(x)=B_{0}+\eta x$, where the gradient $\eta$ is positive during $0\sim t_{a}$ and $3t_{a}+t_{e}\sim 4t_{a}+t_{e}$, while it is negative during $t_{a}\sim 2t_{a}$ and $2t_{a}+t_{e}\sim 3t_{a}+t_{e}$. During the intermediate free flight of duration $t_{e}$, the magnetic field is uniform, and the right path does not accelerate. Such an asymmetrically shaped interferometer can be used to detect gravity-gradients [21, 35, 36]
@@
In this paper, we will focus on the Stern-Gerlach interferometer with the configuration shown in Fig. 1(b), which can be achieved by a system with a spin-1 state and spin-0 state (embedded in an object) in a magnetic field with a constant gradient, see [21]. In particular, the system is prepared in a spin superposition of |$\ket{S=0}$ and |$\ket{S=1}$, say |$\ket{\psi}=1/\sqrt{2}(\ket{S=0}+\ket{S=1})$, while the external magnetic field has a linear spatial distribution in the $x$-axis, i.e. $B(x)=B_{0}+\eta x$, where the magnetic gradient $\eta$ will flip several times to accelerate and decelerate the state |$\ket{S=1}$.
@@
However, some classical noises like RAN and ITN always exist. They will contribute some random phase $\delta\phi$ on $\phi_{\rm diff}$. Although $\mathrm{e}^{\mathrm{i}\delta\phi}$ itself is a pure phase, the ensemble average of such a random phase can lead to a damping factor, known as a dephasing effect. For ease of convenience, $\delta\phi$ is supposed to follow a Gaussian distribution with a mean value of zero and a variance of $\Gamma\equiv\mathbb{E}[(\delta\phi)^{2}]$, then the expectation value of the phase can be computed as 222 The probability density function for the Gaussian distribution is given by $p(x)=\frac{1}{\sqrt{2\pi}\sigma}\mathrm{e}^{-\frac{x^{2}}{2\sigma^{2}}},$ so the expectation value of $\mathbb{E}[\mathrm{e}^{\mathrm{i}\delta\phi}]$ can be directly obtained by computing the integral $\int\frac{1}{\sqrt{2\pi}\sigma}\exp[\mathrm{i}\delta\phi]\exp\left[-\frac{(\delta\phi)^{2}}{2\sigma^{2}}\right]\mathrm{d}(\delta\phi).$
@@
the off-diagonal terms decay exponentially with the damping given by the variance $\Gamma$. Consequently, the expectation value $\text{Tr}(\hat{W}\hat{\rho})$ of any witness operator $\hat{W}$ will also damp by the factor $\mathrm{e}^{-\Gamma}$.
@@ III Dephasing as Linear Response to Inertial Torsion Noise
where $S_{\Theta\Theta}(\omega)$ is the power spectrum density (PSD) of $\Theta(t)$ and the second identity is known as the Wiener-Khinchin theorem  [^3] The strict definition of the PSD of a noise $f_{\rm in}(t)$ is $\lim\limits_{T\to\infty}\frac{|f_{T}(\omega)|^{2}}{T},$ where $f_{T}(\omega)$ is the Fourier transform of $f_{\rm in}(t)$ on the time domain $[0,T]$. We can choose the total experiment time as this $T=4t_{a}+t_{e}$ for a real experiment to calculate the PSD.
The Wiener-Khinchin theorem states that the PSD based on this definition is equivalent to the autocorrelation function of the noise $f_{\rm in}(t)$, that is, $S_{ff}(\omega)=\int\mathbb{E}[f_{\rm in}(t)f_{\rm in}(t+\tau)]\mathrm{e}^{-\mathrm{i}\omega\tau}\mathrm{d}\tau.$ The proof of this theorem can be found in Ref. [46].
. The PSD of $\Theta$ can be measured in the experiment. In this paper, we will consider a specific sources of fluctuation of $\Theta$ (e.g., the collision by gas molecules, which will discussed in the next section).
Generally speaking, there are two dephasing channels in the leading order caused by this noise. First, the noise will contribute a phase along each path according to Feynman’s integral
$\mathrm{e}^{\mathrm{i}\int L_{\rm ITN}[x^{(0)}(t)]/\hbar\mathrm{d}t}.$ (7)
Secondly, the noise will also fluctuate the path of the two arms of the interferometer, so even the ideal phases along the paths will also be disturbed by a phase factor
$\mathrm{e}^{i\int L_{0}[\delta x(t)]/\hbar\mathrm{d}t}.$ (8)
The dephasing effect combining both channels is of higher order and can be neglected.
However, since the noise is assumed to be small, the fluctuations on the trajectories have to be small enough such that the final states of the two arms can meet each other. As a consequence, the second channel of dephasing has to be very small, while the first channel could be significant. Thus, we will focus on the phase fluctuation arising from Eq. (7):
@@
As is shown in Fig.2, $F(\omega)$ tends to a constant $C\sim(\Delta x)^{4}T^{2}\sim 10^{-20}\,\rm m^{4}s^{2}$ in low frequency limit  [^4] According to (12), the factor $\mathrm{e}^{i\omega t}\to 1$ in the low-frequency limit $\omega\to 0$. On the other hand, the superposition size $\Delta x$ is a upper bound for $x_{R}(t)$ while $x_{L}(t)\equiv 0$, so one can obtain $x_{R}^{2}(t)<(\Delta x)^{2}$. Finally, one can get the low-frequency upper bound for $F(\omega)$ from (12) as $F(\omega)<\left|\int(\Delta x)^{2}\mathrm{d}t\right|^{2}=(\Delta x)^{4}T^{2}.$ , where $\Delta x=a_{m}t_{a}^{2}=11.2\,\rm\mu m$ and $T=4t_{a}+t_{e}$ are the superposition size and the total experiment time. On the other hands, $F(\omega)$ decrease as $\omega^{-6}$ in high frequency  [^5] The trajectory $x_{R}(t)$ can be generally expanded as a Taylor series of $t$ with the leading order $x_{R}(t)\sim t$, so $x_{R}(t)\sim t^{2}$. Then the leading order of $F(\omega)$ can be estimated according to (12) as $F(\omega)\sim\left|\int t^{2}\mathrm{e}^{i\omega t}\mathrm{d}t\right|^{2}\sim(\omega^{-3})^{2}=\omega^{-6}.$ So $F(\omega)$ decreases as $\omega^{-6}$ in high-frequency limit. . Combining the low-frequency and high frequency behaviours of $F(\omega)$, one can use the Heaviside step function $\theta(\cdot)$ to approximately describe $F(\omega)$ as
@@ IV Power Spectrum Density of Inertial Torsion Noise
According to the convolution theorem [^6] Suppose there are two time-domain signals $x(t)$ and $y(t)$ with Fourier transform denoted as $X(\omega)$ and $Y(\omega)$, then the convolution theorem states that the Fourier transform of the signal $x(t)y(t)$ is the convolution of $X(\omega)$ and $Y(\omega)$, i.e. $\mathcal{F}[x(t)y(t)]=X(\omega)\ast Y(\omega),$ where the convolution is defined as $X(\omega)\ast Y(\omega)=\int X(u)Y(\omega-u)\mathrm{d}u.$ , the PSD of ITN is the self-convolution of the PSD of the random motion of the suspended apparatus
@@
As is shown in Fig. 1(a), the rotational motion of the experimental box can be modelled as a torsion pendulum. The intrinsic torsion frequency is given by[47]
@@
Considering the gas collision noise caused by gas thermal motion, the dynamical equation of the torsion motion of the experiment box is given by the classical Langevin equation 777 The Langevin equation is a phenomenological motion equation for a classical thermal bath system, see [49]. The physical intuition is that the thermal bath exerts two forces on the system, one of which is a dissipation force proportional to the velocity of the generalized coordinate of the system. At the same time, the other is a random collision force described by a Gaussian process. According to the fluctuation-dissipation theorem, these two forces are represented by the same parameter $\gamma$.  [49]
where $k_{B}$ is the Boltzmann constant and $T=300$ K is the gas temperature outside the experiment box. The input random noise term $\Theta_{\rm in}$ is a dimensionless delta-correlated stationary Gaussian process with zero-mean, satisfying $\mathbb{E}[\Theta_{\rm in}(t)]=0$ at any time $t$, and $\mathbb{E}[\Theta_{\rm in}(t_{1})\Theta_{\rm in}(t_{2})]=\delta(t_{1}-t_{2})$.
According to the fluctuation-dissipation theorem, the damping rate $\gamma_{\rm rot}$ describes both the dissipation effect and the random force caused by the collision from the ambient thermal gas molecules, which is given by [50, 51]
where $P_{\rm gas}$ is the pressure of gas and $m_{\rm gas}$ is the mass of the gas molecules. Since $\gamma_{\rm rot}$ is proportional to $P_{\rm gas}$, it can vary depending on the gas pressure outside the box. Apart from $P_{\rm gas}$, all other factors contribute a factor around $10^{-4}\sim 10^{-3}$, so the value of the damping rate $\gamma_{\rm rot}$ can be estimated as
$\gamma_{\rm rot}/[\rm Hz]\sim 10^{-4}P_{\rm gas}/[\rm Pa].$ (20)
@@
$S_{\Theta\Theta}(\omega)=\frac{2\gamma_{\rm rot}k_{B}T/I}{(\Omega_{\rm rot}^{2}-\omega^{2})^{2}+\gamma_{\rm rot}^{2}\omega^{2}}.$ (21)
Figure 3: Monte Carlo simulation for the motion of the box. (a) Simulated time-trace of Eq. (18). (b) The PSD $S_{\Theta\Theta}$ in the lower plot is computed by FFT. Theoretically the PSD is a Lorentzian distribution which fits the simulation result well. The intrinsic frequency $\Omega_{\rm rot}$ is chosen as $2\pi\times 1$ Hz, the damping rate $\gamma_{\rm rot}$ is chosen as $10^{-10}$ Hz, the temperature is set as 300 K.
@@
Since the PSD of ITN is the self-convolution of $S_{\dot{\Theta}\dot{\Theta}}(\omega)$ according to Eq. (15), one can compute the PSD of the ITN driven by thermal gas collisions with the PSD given by Eq. (21) as
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)=\bar{S}\frac{4\omega^{4}+4(\gamma_{\rm rot}^{2}-3\Omega_{\rm rot}^{2})\omega^{2}+16\Omega_{\rm rot}^{4}}{(\omega^{2}+\gamma_{\rm rot}^{2})(4\gamma_{\rm rot}^{2}\omega^{2}+(\omega^{2}-4\Omega_{\rm rot}^{2})^{2})},$ (22)
where the prefactor is given by
$\bar{S}=\frac{4\pi\gamma_{\rm rot}(k_{B}T)^{2}}{I^{2}}.$ (23)
@@
Figure 4: PSD of ITN $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ and angular motion PSD $S_{\Theta\Theta}(\omega)$. The parameters are chosen as $\Omega_{\rm rot}=2\pi\times 1$ Hz, $\gamma_{\rm rot}=10^{-10}$ Hz, I=10 $\rm kg\cdot m^{2}$ and T=300 K. $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is numerically orders of magnitude smaller than the value of $S_{\Theta\Theta}(\omega)$ because it has an additional $k_{B}T\sim 10^{-21}$ J factor as can be seen by comparing Eqs. (21) and  (22) (the two PSDs have the units shown on the right and left axis, respectively). The peak frequency of $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is doubled compared to $S_{\Theta\Theta}(\omega)$ as expected when computing the square of the noise 9. The dashed green and purple lines show that the theoretical expressions in Eqs. (21) and  (22) fit well the simulated PSDs.
@@
The detailed mathematical steps are summarized in Appendix B. Fig. 4 shows the analytical result and the numerical simulation of both $S_{\Theta\Theta}(\omega)$ and $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$. As is shown, the magnitude of ITN is about $10^{-20}$ level smaller than $S_{\Theta\Theta}$. This is because $S_{\Theta\Theta}(\omega)$ is proportional to $k_{B}T$ but $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is proportional to $(k_{B}T)^{2}$ which contributes an additional factor $k_{B}T\sim 10^{-21}$ J.
@@
The asymptotic behaviour of the PSD of ITN is also different from the $S_{\Theta\Theta}(\omega)$. In low frequency limit, $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ behaves like $1/(\omega^{2}+\gamma_{\rm rot}^{2})$ and tends to a constant $4\pi(k_{B}T/I)^{2}/\gamma_{\rm rot}$. In high frequency region, $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ decreases as $\omega^{-2}$ according to the analytic result (22).

### v2

@@ Abstract
Matter-wave interferometry is susceptible to non-inertial noise sources, which can induce dephasing and a resulting loss of interferometric visibility. Here, we focus on inertial torsion noise (ITN), which arises from the rotational motion of the experimental apparatus suspended by a thin wire and subject to random external torques. We provide analytical expressions for the ITN noise starting from generalized Langevin equations describing the experimental box which can then be used together with the transfer function to obtain the dephasing factor. We verify the theoretical modelling and the validity of the approximations using Monte Carlo simulations, obtaining good agreement between theory and numerics. As an application, we estimate the size of the effects for the next generation of interferometry experiments with femtogram particles, which could be used as the building block for entanglement-based tests of the quantum nature of gravity. We find that the ambient gas is a weak source of ITN, posing mild restrictions on the ambient pressure and temperature. We discuss the general ITN constraints by assuming a Langevin equation parameterized by three phenomenological parameters.
@@ I Introduction
Matter-wave interferometry has many salient applications for gravitational physics with devices spanning gravimeters [1], gradiometers [2, 3], accelerometers [4] and gyroscopes [5]. They can also be used for fundamental physics, such as testing the equivalence principle [6, 7, 8] and the quantum gravity induced entangled of masses (QGEM) to test the quantum nature of gravity in a lab [9, 10], see also [11, 12, 13, 14]. The paper [14] provides the protocol to test the spin-2 nature of the graviton in an analogue of the light-bending experiment, see also [13]. One can also probe the nature of massive graviton [15] and non-local gravitational interaction [16] motivated by string theory. Furthermore, some groups even consider building gravitational wave observatories based on matter-wave interferometry such as the matter-wave laser interferometer gravitation antenna (MIGA) [17, 18], the matter-wave atomic gradiometer interferometric sensor (MAGIS-100) [19, 20], as well as the mesoscopic interference for metric and curvature (MIMAC) scheme [21].
Future matter-wave interferometry aims to exploit the regime of large masses, large superposition sizes, and long coherence times, allowing the probe of exquisitely small experimental signals. For example, the QGEM experiment would ideally require a test mass of $\sim 10^{-15}$ kg, a superposition size of $\sim 100\,\rm\mu m$, and a coherence time of $\sim$1 s [9, 8, 22, 23, 24]. One of the most promising setups towards this kind of experiments is the adaptation of the Stern-Gerlach interferometer (SGI) to nanoparticles [25]. SGIs based on atom chips [26] have already achieved the superposition size and coherence time of 3.93 $\rm\mu m$ and 21.45 ms for the half-loop configuration [27], respectively, and $0.38\,\rm\mu m$ and 7 ms for the full-loop configuration [28], respectively. The next generation of SGIs is currently under theoretical and numerical investigation [29, 30, 31, 32, 33].
@@
This paper will study the dephasing caused by the residual rotational or inertial torsion noise (ITN) for an asymmetric nanoparticle matter-wave interferometer which is sensitive to gravity-gradients [21, 35, 36]. ITN arises naturally in any setup whenever the experimental apparatus is subject to random torques, placing it in non-inertial rotational motion. As we will see, ITN can induce relative random phases in an interferometric experiment, resulting in a loss of interferometric contrast. Here, we will be primarily interested in understanding ITN, and we will focus on a simple single-stage suspension forming a torsion pendulum, i.e., matter-wave interferometry performed inside a hanging box. Analogous configurations of the experimental apparatus have been investigated previously in gravitational wave observatories like LIGO [40] and VIRGO [41]. More advanced setups could employ additional structures like the inverted pendulum [42] and the Roberts linkage [43].
In Sec. II, we first introduce the concept of ITN, illustrating it for matter-wave interferometry that can be modelled using qubits. In Sec. III, we generalize the analysis using linear response theory, providing the transfer function and its relation to the dephasing factor. In Sec. IV, we investigate ITN caused by ambient gas collisions on the experimental box starting from a generalized classical Langevin equation. We will compute the power spectrum density (PSD) of the ITN using the convolution theorem and verify the validity of the approximations using Monte Carlo simulations. In Sec. V, we obtain the resulting constraints on the ambient pressure and temperature and general constraints on ITN by parametrizing a generic Langevin equation modelling the experimental box with three phenomenological parameters in the dynamical equation of the experimental box. In Sec. VI, we conclude with a summary of the obtained results. In Appendix A, we construct the ITN Lagrangian starting from Fermi normal coordinates and transforming to a rotating reference frame. In Appendix B, we provide for completeness and also the complete derivation of the ITN PSD using complex analysis. In Appendix C, we analyze the ITN caused by the ambient thermal gas collision.
II Concept of Inertial Torsion Noise and Dephasing
@@
Figure 1: (a) Experimental scheme and illustration of inertial torsion noise (ITN). The matter-wave particle (light and dark blue circles) is placed inside the experimental box (blue box), which is suspended by a thin wire (black vertical line). The box can rotate around the z-axis, while the interferometric protocol is performed along the horizontal x-axis. ITN generates random torques on the experimental box (blue rectangle), placing it in non-inertial rotational motion. The random non-inertial rotational motion induces random relative phases, which can lead to dephasing and the loss of interferometric contrast. (b) The structure of a Stern-Gerlach interferometer with spin-1 [25]. The left path keeps the same position, while the right path accelerates and decelerates at a constant rate, resulting in a superposition size $\Delta x$. Such motion is achieved using a constant magnetic field gradient, i.e. $B(x)=B_{0}+\eta x$, where the gradient $\eta$ is positive during $0\sim t_{a}$ and $3t_{a}+t_{e}\sim 4t_{a}+t_{e}$, while it is negative during $t_{a}\sim 2t_{a}$ and $2t_{a}+t_{e}\sim 3t_{a}+t_{e}$. During the intermediate free flight of duration $t_{e}$, the magnetic field is uniform, and the right path does not accelerate. Such an asymmetrically shaped interferometer can be used to detect gravity-gradients [21, 35, 36]
@@
In this paper, we will focus on the Stern-Gerlach interferometer with the configuration shown in Fig. 1(b), which can be achieved by a system with a spin-1 state and spin-0 state (embedded in an object) in a magnetic field with a constant gradient, see [21]. In particular, the system is prepared in a spin superposition of |$\ket{S=0}$ and |$\ket{S=1}$, say |$\ket{\psi}=1/\sqrt{2}(\ket{S=0}+\ket{S=1})$. At the same time, the external magnetic field has a linear spatial distribution in the $x$-axis, i.e. $B(x)=B_{0}+\eta x$, where the magnetic gradient $\eta$ will flip several times to accelerate and decelerate the state |$\ket{S=1}$.
@@
However, some classical noises like RAN and ITN always exist. They will contribute some random phase $\delta\phi$ on $\phi_{\rm diff}$. Although $\mathrm{e}^{\mathrm{i}\delta\phi}$ itself is a pure phase, the ensemble average of such a random phase can lead to a damping factor, known as a dephasing effect. For ease of convenience, $\delta\phi$ is supposed to follow a Gaussian distribution with a mean value of zero and a variance of $\Gamma\equiv\mathbb{E}[(\delta\phi)^{2}]$/2, then the expectation value of the phase can be computed as
@@
the off-diagonal terms decay exponentially with the damping given by the variance $\Gamma$. Consequently, the expectation value $\text{Tr}(\hat{W}\hat{\rho})$ of any witness operator $\hat{W}$ will also decay with respect to the factor $\mathrm{e}^{-\Gamma}$.
For example, one can consider two applications of matter-wave interferometers for gravity experiments.
  • The first application is the gravimeter based on a NV center [44, 46]. The gravitational acceleration $g$ is proportional to the differential phase as $\phi_{\rm diff}=16\pi mg\Delta z/(\hbar\omega_{0})$, where $\Delta z$ is the superposition size along the z-axis, and $\omega_{0}\sim 100$ kHz is the trapping frequency. Then the phase fluctuation can cause a sensitivity loss on the measurement result of $g$, then one can obtain a constrain for $\Gamma$ by the variance $\sigma_{g}$ of $\delta g$ as $\sqrt{\Gamma}=16\pi m\sigma_{g}\Delta z/(\hbar\omega_{0})$, as long as $\delta\phi_{\rm diff}$ and $\delta g$ follow Gaussian distribution. Choosing the values as $m\sim 10^{-16}$ kg, $\Delta z\sim 10^{-8}$ m [47], then one may obtain a threshold as $\sqrt{\Gamma}<\sigma_{g}10^{7}\,\rm{s^{2}/m}$. A common performance of gravimeters is $\sigma_{g}=10^{-9}\sim 10^{-10}g$ [47], then one can choose the threshold of $\Gamma$ as $10^{-6}$.
  • Another example is the QGEM experiment [38, 35], which uses two interferometers coupled by gravity to investigate gravity-induced quantum entanglement. The witness $\hat{W}$ is proposed as the positive partial transpose (PPT) witness, which, in the case of two qubits, provides a sufficient and necessary condition for entanglement based on the Peres-Horodecki criterion [48, 49, 50], which requires the expectation value of the witness satisfying $\braket{\hat{W}}<0$. As is calculated in [38], the expectation value of the witness under the dephasing effect is $\braket{\hat{W}}=(1-\mathrm{e}^{-2\Gamma})/4-\mathrm{e}^{-\Gamma/2}\sin\phi_{g}$, then the constrain on $\Gamma$ is that $\Gamma/2<\mathrm{e}^{-\Gamma/2}\sin\phi_{g}\approx\phi_{g}$. The value of $\phi_{g}$ has been estimated as $\phi_{g}\approx 0.015$ [51, 35], then we will choose a threshold $\Gamma=0.01$ for further discussions on the values of parameters in this paper.
@@ III Dephasing as Linear Response to Inertial Torsion Noise
where $S_{\Theta\Theta}(\omega)$ is the power spectrum density (PSD) of $\Theta(t)$ and the second identity is known as the Wiener-Khinchin theorem. The PSD of $\Theta$ can be measured in the experiment. In this paper, we will consider a specific source of fluctuation of $\Theta$ (e.g., the collision by gas molecules, which will discussed in the next section).
As is proved in [52, 53], the phase fluctuations due to noises are determined by the path integral of the corresponding Lagrangian of the noise along the unperturbed classical trajectories. Thus, the phase fluctuation is
@@
As is shown in Fig.2, $F(\omega)$ tends to a constant $C\sim(\Delta x)^{4}T^{2}\sim 10^{-20}\,\rm m^{4}s^{2}$ in low frequency limit where $\Delta x=a_{m}t_{a}^{2}=11.2\,\rm\mu m$ and $T=4t_{a}+t_{e}$ are the superposition size and the total experiment time. It is because the factor $\mathrm{e}^{i\omega t}\to 1$ in the low-frequency limit $\omega\to 0$, and the superposition size $\Delta x$ is a upper bound for $x_{R}(t)$ when $x_{L}(t)\equiv 0$, so $F(\omega)$ in (10) can be approximated in low-frequency region as
$F(\omega)\sim\left|\int(\Delta x)^{2}\mathrm{d}t\right|^{2}=(\Delta x)^{4}T^{2}.$ (12)
At the high frequency limit, $F(\omega)$ decrease as $\omega^{-6}$. In particular, the trajectory $x_{R}(t)$ can be generally expanded as a Taylor series of $t$ with the leading order $x_{R}(t)\sim t$, so $x^{2}_{R}(t)\sim t^{2}$, then the leading order of $F(\omega)$ can be estimated according to (10) as
$F(\omega)\sim\left|\int t^{2}\mathrm{e}^{i\omega t}\mathrm{d}t\right|^{2}\sim(\omega^{-3})^{2}=\omega^{-6}.$ (13)
Combining the low-frequency and high-frequency behaviours of $F(\omega)$, one can use the Heaviside step function $\theta(\cdot)$ to approximately describe $F(\omega)$ as
@@ IV Power Spectrum Density of Inertial Torsion Noise
According to the convolution theorem, the PSD of ITN is the self-convolution of the PSD of the random motion of the suspended apparatus
@@
As is shown in Fig. 1(a), the rotational motion of the experimental box can be modelled as a torsion pendulum, which follows a generalized Langevin equation
$\ddot{\Theta}=-\Omega_{\rm rot}^{2}\Theta-\gamma\dot{\Theta}+\sqrt{A}\Theta_{\rm in}(t).$ (17)
The input random noise term $\Theta_{\rm in}(t)$ is a unit delta-correlated stationary Gaussian process with zero-mean, satisfying $\mathbb{E}[\Theta_{\rm in}(t)]=0$ at any time $t$, and $\mathbb{E}[\Theta_{\rm in}(t_{1})\Theta_{\rm in}(t_{2})]=\delta(t_{1}-t_{2})$. The amplitude of the external noise $\sqrt{A}$  [^1] Note that $A$ is of the unit $\rm Hz^{4}/Hz$. It is because the term $\sqrt{A}\Theta_{\rm in}(t)$ has a dimension $\rm[T^{-2}]$, where $\Theta_{\rm in}(t)$ has a dimension $\rm[T^{-1/2}]$. and the dissipation rate $\gamma$ are determined by the experiment. The intrinsic torsion frequency is given by [54]
@@
$S_{\Theta\Theta}(\omega)=\frac{A}{(\Omega_{\rm rot}^{2}-\omega^{2})^{2}+\gamma^{2}\omega^{2}}.$ (19)
Figure 3: Monte Carlo simulation for the motion of the box. (a) Simulated time-trace of Eq. (39). (b) The PSD $S_{\Theta\Theta}(\omega)$ in the lower plot is computed by FFT. Theoretically the PSD is a Lorentzian distribution which fits the simulation result well. The intrinsic frequency $\Omega_{\rm rot}$ is chosen as $2\pi\times 1$ Hz, the damping rate $\gamma$ is chosen as $10^{-10}$ Hz and the noise amplitude $\sqrt{A}$ is chosen as $10^{-3}\,\rm Hz^{2}/\sqrt{Hz}$
@@
Since the PSD of ITN is the self-convolution of $S_{\dot{\Theta}\dot{\Theta}}(\omega)$ according to Eq. (15), one can compute the PSD of the ITN driven by the external noise with the PSD given by Eq. (19) as
$S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)=A^{2}\frac{4\omega^{4}+4(\gamma^{2}-3\Omega_{\rm rot}^{2})\omega^{2}+16\Omega_{\rm rot}^{4}}{\gamma(\omega^{2}+\gamma^{2})(4\gamma^{2}\omega^{2}+(\omega^{2}-4\Omega_{\rm rot}^{2})^{2})}.$ (20)
@@
Figure 4: PSD of ITN $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ and angular motion PSD $S_{\Theta\Theta}(\omega)$. The parameters are chosen as $\Omega_{\rm rot}=2\pi\times 1$ Hz, $\gamma=10^{-10}$ Hz, I=10 $\rm kg\cdot m^{2}$ and $\sqrt{A}=10^{-3}\,\rm Hz^{2}/\sqrt{Hz}$. . The peak frequency of $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ is doubled compared to $S_{\Theta\Theta}(\omega)$ as expected when computing the square of the noise 3. The dashed green and purple lines show that the theoretical expressions in Eqs. (19) and  (20) fit well the simulated PSDs.
@@
The detailed mathematical steps are summarized in Appendix B. Fig. 4 shows the analytical result and the numerical simulation of both $S_{\Theta\Theta}(\omega)$ and $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$.
@@
The asymptotic behaviour of the PSD of ITN is different from the $S_{\Theta\Theta}(\omega)$. In low frequency limit, $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ behaves like $1/(\omega^{2}+\gamma_{\rm rot}^{2})$ and tends to a constant $A^{2}/\gamma$. In high frequency region, $S_{\dot{\Theta}^{2}\dot{\Theta}^{2}}(\omega)$ decreases as $\omega^{-2}$ according to the analytic result (20).